# Tales of stellar and binary co-evolution, told by stellar oscillations and age estimates in main-sequence and red-giant stars

### v1

@@ 3.2 Binary fraction evolution for solar-like stars into the RC
Solar-like oscillating dwarfs constitute the progenitor sample for the giant stars that follow the evolutionary channel leading into degenerate He-core ignition. We, therefore, use the sample of 620 oscillating solar-like MS and subgiant (SG) stars observed by Kepler (Chaplin et al., 2011; Mathur et al., 2022) to estimate the main-sequence baseline binary fraction.
@@
Using the Union Binary Flag from Godoy-Rivera et al. (2025), we identify 195 binary candidates in this sample, yielding a binary fraction of 31.5%. This value is in good agreement with the binary fraction of 35% reported by Badenes et al. (2018) for main-sequence stars in the sample of APOGEE, showing that for this sample, the possible selection bias potentially introduced by basing the analysis on solar-like oscillating stars is small for the reference sample. However, both values are well below the expected 50%, suggested by large-sample statistics. For Badenes et al. (2018), this discrepancy can be explained by the nature of APOGEE spectroscopy, where some binaries remain undetected due to the suboptimal sampling of the observations or inclination effects. The binary indicators in the Union Binary Flag are dominated by Gaia RUWE measurements (accounting for 76.5% of binary candidate indicators). As mentioned before, 40% of confirmed SB9 binaries have RUWE 1.4 (Beck et al., 2024), which suggests that a significant fraction of binaries in our sample might remain undetected due to weak photocentric motion.
@@
For the computation of the binary rate on the main sequence, we should consider that this sample could be biased against binarity. From the sample of Mathur et al. (2022) we find that about 50% of all dwarfs observed by Kepler in the one-minute (short cadence) sampling show solar-like oscillations. Mathur et al. (2019) argued that the non-detection of mode is probably caused by high levels of activity, low metallicity, or tidal interaction, which reduce the oscillation amplitudes below the detection limit. The latter assumption was motivated by the finding that red giants in binary systems show suppressed oscillation, once the fractional radius of both components fills more than 20% of the semi-major axis, a, 20%, and the tidal spin up activates a stellar dynamo (Gaulme et al., 2014). Such a fractional radius coincides with the threshold of tidal interaction through the equilibrium tide for the circularization and synchronization (Beck et al., 2018b). However, Beck et al. (2024) showed that solar-like stars in strongly interacting binary systems that are synchronized and circularized by the dynamical tide ( 8 days, Zahn, 1989; Raghavan et al., 2010) can exhibit solar-like oscillations even at high levels of photospheric activity. Among the 663 MS and SG stars that were observed in short cadence and did not show a significant excess of oscillation, we find 270 binary candidates from the binary union flag, which corresponds to a binary fraction of 40.7%, which suggests that binarity might have a small effect but is not the dominant mechanism to reduce the amplitude of solar-like oscillations in MS stars. Therefore, we consider this base value provides a good reference for the binary fraction on the main sequence.
@@
For stars with masses below 1.8 M⊙, we find 438 binary candidates with RGB primaries and 200 with primaries in the RC, corresponding to binary fractions of 8.2% for the H-shell and 5.2% for the He-core burning stars that exhibit solar-like oscillations (Table 3 and Fig. 5). From these numbers, we estimate a binary attrition from the main sequence to the RGB phase to be 69% (Table 6), which is in good agreement with the reduction found in the APOGEE sample (Badenes et al., 2018). The binary depletion is even larger between the main-sequence and the red-clump phase with 84%.
With a maximum change of the stellar radius of nearly two orders of magnitude, the RGB is the phase with the most significant variations and differences in stellar radius. Therefore, a significant reduction along the RGB evolution itself can be expected. To test this variation, we use the peak frequency of the oscillatory power excess, , to split the RGB into two subsamples and evaluate their binary fractions separately. For the low-luminosity RGB (300 30 Hz, which correspond to approximately to a range in radius of 3 R/R⊙ 10) we find a binary fraction of 9.7%. For the high-luminosity RGB (30 0.1 Hz, 10 R/R⊙ 100) we find a binary fraction of 6.0%. This corresponds to an attrition of 38% from the low- to the high-luminosity RGB, of 81% with respect to the MS. For the detailed overview of the values, we refer the reader to Fig. 3, as well as Tables 5 and 6.
@@
We applied the test for independence to evaluate whether the observed differences in binary fractions between evolutionary states are statistically significant, under the null hypothesis that binarity is independent of stellar evolutionary phase. We adopt the traditional threshold-based classification of -values, whereby results with 0.001 are deemed highly significant, indicating strong statistical evidence against the null hypothesis. When -values fall in the range 0.001 0.05, the variation is deemed marginally significant, suggesting a moderate likelihood that the observed difference is not purely due to statistical fluctuations. Conversely, if 0.05, we classify the variation as insignificant. The significance test, whose results are reported in Table 6 and depicted in Fig. 3, shows that the binary attrition from the main sequence to any state of the giant branch and the RGB to the RC are significant. It is worth highlighting that the 38% reduction between the low- and high-luminosity regimes of the RGB evolution is highly significant. The only variation not found to be significant in the low-mass sample is the reduction of 15% from the high-luminosity RGB into the RC phase. To determine whether the small variation between the high-luminosity RGB and the RC is real, a larger sample would be required.
Finally, we note that two important potential biases exist for this analysis. First, the selection criterion on solar-like oscillations introduces a bias. As mentioned before, it was shown that solar-like oscillations are suppressed in giants where the combined fractional stellar radii exceed 20% of the system’s semi-major axis. Consequently, stars starting to fill their Roche-lobe substantially will also be excluded from the sample of oscillating giants. However, once this phase has started, the time scales of a few 100 kiloyears, leading into a potentially destructive evolution and white-dwarf phase, are short. Given the duration of the RGB phase, this dilution effect does not significantly skew the sample. Second, stellar oscillations are also detectable from RV variations. However, only on the super-luminous RGB ( 0.1 Hz, or 100 R/R⊙ 200), the RV amplitude of stellar oscillations becomes substantial enough to reach or exceed the RV threshold of 1.4 km/s (Katz et al., 2019) for faint stars of the Gaia RVS instrument ( 12 mag). As this regime is not probed by the range in APOKASC 3, possible false positive binary solutions produced by intrinsic stellar RV variations do not bias this comparison. We, therefore, assume the completeness constant throughout the studied evolutionary states.
@@ 3.3 Evolution of the binary fraction for 2RC giants
The progenitors of secondary clump stars are early F- and A-type stars, which exhibit a different type of pulsations ( Dor and Scuti, Aerts et al., 2010, and references therein). Therefore, these are not represented in the main-sequence sample of Mathur et al. (2022). We measure the binary fraction for the non-degenerate evolution starting from the low-luminosity RGB.
Table 5 and Fig. 2 show that such relatively massive stars represent only a small fraction of the full APOKASC 3 sample. Consequently, their population and binary counts suffer from small-sample statistics, and all differences found between the subsamples were deemed non-significant. The only marginally significant difference that we found from the numbers in Table 6 was that the binary fraction of the full sample of non-degenerate RGB stars (9.6%) is 17% richer in binaries than the full sample of degenerate RGB (8.2%). This increased richness aligns with the differences in the binary fraction of their main-sequence progenitors, as observed in large binary statistics (Moe & Di Stefano, 2017; Offner et al., 2023).
4 Co-evolution of the primary mass and the orbit
The key parameter determining the evolution of a star is its initial mass (Serenelli et al., 2021, and references therein). It defines the stellar structure and timescales of the evolution. In a binary system, the total mass of the components in solar units (), together with the semi-major axis, (in AU) or orbital period in years, defines the system’s configuration through the third law of Kepler (1619),
@@ 4 Co-evolution of the primary mass and the orbit
Tidal interactions gain importance when the radius of one of the stellar components grows to a significant fraction of the semi-major axis. The redistribution of angular momentum and dissipation of kinetic and potential energy from tides into heat drives the changes in the orbits and the spins. This process leads the system toward a state of minimum energy, where the orbit is circularized, and the stellar spins become aligned and synchronized (Zahn, 1977; Ogilvie & Lin, 2007; Mathis, 2015). Tidal interaction occurs in waves as a dynamical tide or as a hydrostatic adjustment in the equilibrium tide, whereby the stellar structure, which depends primarily on the stellar mass, determines the dominant tidal mechanism. The various manifestations of tides are intensively studied through the distributions of orbital parameters in large samples from open clusters or from the Gaia mission, providing valuable constraints on tidal dissipation and orbital evolution (e.g. Mirouh et al., 2023; Dewberry & Wu, 2025).
Mass loss during stellar evolution, such as stellar winds, also affects orbital evolution by removing mass and angular momentum from the system (Miglio et al., 2012). Following Kepler’s third law (Eq. 1), this leads to an increase of the orbital period (Zahn & Bouchet, 1989; Soberman et al., 1997) and also to the spin-down of the mass-loosing star (e.g. Skumanich, 1972).
@@
If the radius of one of the binary components exceeds the orbit of the companion, the system will enter the short-lived, most likely destructive common-envelope (CE) phase, which would end the red-giant phase of the star abruptly with the CE ejection, exposing the ”naked” He-core as a white dwarf (Ivanova et al., 2013). Due to the hydrodynamical complexity and their short time scales, this phase is difficult to model and poorly understood (Han et al., 2002). However, several parameters are identified to play a significant role in determining the fate of the binary system. The primary determining factor is the semi-major axis relative to the maximum stellar radius of the components, as this governs both tidal interaction strength and the Roche-lobe radius. Furthermore, the mass ratio between the two stars, which influences the difference in their evolutionary timescales (e.g. Miglio et al., 2014; Beck et al., 2018a; Grossmann et al., 2025)
@@
Since stellar ages are typically inferred from model fitting, based on single-star evolutionary models assuming steady mass loss (e.g. Pinsonneault et al., 2025), episodic or interaction-induced mass loss in binary systems will significantly over- or underestimate the ages of the post-interaction stellar products.
@@ 4.1 Mass distributions separated by evolutionary state
To isolate evolutionary effects in the binary sample, we construct reference mass distributions for the entire sample, separated by evolutionary state. Since the mass distributions differ between the APOKASC 3 and APO-K2 catalogs, we analyze these two catalogs separately to avoid potential systematic biases in the samples skewing the result. The resulting KDE, shown in Fig. 4, shows that the mass distribution among the full samples is similar. Furthermore, Fig. 4 presents KDEs for the mass distributions for samples selected from various binary indicators. In the RGB phase, the mass distribution of all binary markers closely resembles that of the full sample, with only minor deviations.
@@
We suggest that the upward shift of the median mass in the binary sample is caused by the removal of lower-mass systems that arise from disruptive and merging events along the RGB, as shorter-period binaries undergo a CE phase, mostly leading to merger or envelope stripping of the RGB primary, before such stars can reach the tip of the giant branch and continue into the RC phase. This mechanism preferentially removes systems with and days and introduces a stronger mass dependence on the binary fraction than for the main-sequence progenitors. Because the orbital period is the main parameter determining whether a system undergoes the CE phase, we expect a mass–period relation in the He-core-burning phase.
@@ 4.2 Mass dependency of the orbital parameters
For intermediate periods (8 Porb [d] 104), orbital eccentricities of MS binaries follow a Maxwellian ”thermal” distribution (Zahn, 1989; Raghavan et al., 2010; Moe & Di Stefano, 2017), indicating that the probability density increases linearly with , i.e., systems with higher eccentricities are more common than low or circularized eccentricities which is still present in the eccentricities of main-sequence stars in Fig. 5.
The combined large sample of binary stars with orbital solutions from the TBO with seismically inferred evolutionary states and masses from the APOKASC 3 and APO-K2 allows us to populate the eccentricity-period (e-P) plane (Fig. 5) and study the binary evolution as a function of stellar evolution (Fig. 6). We find a significant reduction in the ranges of orbital period and eccentricity from the main sequence into the He-core-burning phase. The short-periodic end of the period distribution shifts from a few days for MS stars to about 20 days for RGB. For stars in the He-core burning phase, the short periodic end is about 200 and 500 days for 2RC and RC stars. This picture is consistent with the findings of Dewberry & Wu (2025), who quantified comparable trends from a large-sample study of low-mass stellar binaries in the Gaia DR3 catalog.
@@ 4.3 Post-common envelope systems
We find numerous short periodic systems among the systems shown in the e-P plane in Fig. 5. Such systems are most likely artifacts of longer-periodic systems in the TBO. While red giants could exist in short-period systems, it is unlikely that such primaries would exhibit solar-like oscillations due to the effects of tidal interactions. We, therefore, excluded systems with Porb,TBO 10 days from the analysis but show them in Fig. 5.
@@
Only 7.5% of all giants with enhanced surface rotation are found to be in binary systems. Although this is slightly above the overall binary rate for RC stars, it is notable, as Patton et al. (2024) reported, based on all available binary indicators from Gaia DR3 and APOGEE RVs, a binary fraction of 732% for the entire sample of giants with enhanced surface rotation. This is significantly higher than the expected binary fraction of low-mass main-sequence stars, indicating that these are the products of mergers and disruptive events.

### v2

@@ 3.2 Binary fraction evolution for solar-like stars into the RC
Solar-like oscillating dwarfs constitute the progenitor sample for the giant stars that follow the evolutionary channel leading into degenerate He-core ignition. We therefore used the sample of 620 oscillating solar-like MS and subgiant (SG) stars observed by Kepler (Chaplin et al., 2011; Mathur et al., 2022) to estimate the MS baseline binary fraction.
@@
Using the Union Binary Flag from Godoy-Rivera et al. (2025), we identify 195 binary candidates in this sample, yielding a binary fraction of 31.5%. This value is in good agreement with the binary fraction of 35% reported by Badenes et al. (2018) for MS stars in the sample of APOGEE, showing that for this sample, the possible selection bias potentially introduced by basing the analysis on solar-like oscillating stars is small for the reference sample. However, both values are well below the expected 50%, suggested by large-sample statistics. For Badenes et al. (2018), this discrepancy can be explained by the nature of APOGEE spectroscopy, where some binaries remain undetected due to the suboptimal sampling of the observations or inclination effects. The binary indicators in the Union Binary Flag are dominated by Gaia RUWE measurements (accounting for 76.5% of binary candidate indicators). As mentioned before, 40% of confirmed SB9 binaries have RUWE 1.4 (Beck et al., 2024), which suggests that a significant fraction of binaries in our sample might remain undetected due to weak photocentric motion.
@@
For the computation of the binary rate on the MS, we should consider that this sample could be biased against binarity. From the sample of Mathur et al. (2022) we find that about 50% of all dwarfs observed by Kepler in the one-minute (short cadence) sampling show solar-like oscillations. Mathur et al. (2019) argued that the non-detection of mode is probably caused by high levels of activity, low metallicity, or tidal interaction, which reduce the oscillation amplitudes below the detection limit. The latter assumption was motivated by the finding that red giants in binary systems show suppressed oscillation, once the fractional radius of both components fills more than 20% of the semimajor axis, a, 20%, and the tidal spin up activates a stellar dynamo (Gaulme et al., 2014). Such a fractional radius coincides with the threshold of tidal interaction through the equilibrium tide for the circularization and synchronization (Beck et al., 2018b). However, Beck et al. (2024) showed that solar-like stars in strongly interacting binary systems that are synchronized and circularized by the dynamical tide ( 8 days, Zahn, 1989; Raghavan et al., 2010) can exhibit solar-like oscillations even at high levels of photospheric activity. Among the 663 MS and SG stars that were observed in short cadence and did not show a significant excess of oscillation, we find 270 binary candidates from the binary union flag, which corresponds to a binary fraction of 40.7%, which suggests that binarity might have a small effect but is not the dominant mechanism to reduce the amplitude of solar-like oscillations in MS stars. Therefore, we consider this base value provides a good reference for the binary fraction on the MS.
@@
For stars with masses below 1.8 M⊙, we find 438 binary candidates with RGB primaries and 200 with primaries in the RC, corresponding to binary fractions of 8.2% for the H-shell and 5.2% for the He-core burning stars that exhibit solar-like oscillations (Table 3 and Fig. 5). From these numbers, we estimate a binary attrition from the MS to the RGB phase to be 69% (Table 6), which is in good agreement with the reduction found in the APOGEE sample (Badenes et al., 2018). The binary depletion is even larger between the MS and the RC phase with 84%.
With a maximum change of the stellar radius of nearly two orders of magnitude, the RGB is the phase with the most significant variations and differences in stellar radius. Therefore, a significant reduction along the RGB evolution itself can be expected. To test this variation, we used the peak frequency of the oscillatory power excess, , to split the RGB into two subsamples and evaluate their binary fractions separately. For the low-luminosity RGB (300 30 Hz, which correspond to approximately to a range in radius of 3 R/R⊙ 10) we find a binary fraction of 9.7%. For the high-luminosity RGB (30 0.1 Hz, 10 R/R⊙ 100) we find a binary fraction of 6.0%. This corresponds to an attrition of 38% from the low- to the high-luminosity RGB, of 81% with respect to the MS. For the detailed overview of the values, we refer the reader to Fig. 3, as well as Tables 5 and 6.
@@
We applied the test for independence to evaluate whether the observed differences in binary fractions between evolutionary states are statistically significant, under the null hypothesis that binarity is independent of stellar evolutionary phase. We adopt the traditional threshold-based classification of values, whereby results with 0.001 are deemed highly significant, indicating strong statistical evidence against the null hypothesis. When values fall in the range 0.001 0.05, the variation is deemed marginally significant, suggesting a moderate likelihood that the observed difference is not purely due to statistical fluctuations. Conversely, if 0.05, we classify the variation as insignificant. The significance test, whose results are reported in Table 6 and depicted in Fig. 3, shows that the binary attrition from the MS to any state of the giant branch and the RGB to the RC are significant. It is worth highlighting that the 38% reduction between the low- and high-luminosity regimes of the RGB evolution is highly significant. The only variation not found to be significant in the low-mass sample is the reduction of 15% from the high-luminosity RGB into the RC phase. To determine whether the small variation between the high-luminosity RGB and the RC is real, a larger sample would be required.
Finally, we note that two important potential biases exist for this analysis. First, the selection criterion on solar-like oscillations introduces a bias. As mentioned before, it was shown that solar-like oscillations are suppressed in giants where the combined fractional stellar radii exceed 20% of the system’s semimajor axis. Consequently, stars starting to fill their Roche-lobe substantially will also be excluded from the sample of oscillating giants. However, once this phase has started, the timescales of a few hundred kiloyears, leading into a potentially destructive evolution and white-dwarf phase, are short. Given the duration of the RGB phase, this dilution effect does not significantly skew the sample. Second, stellar oscillations are also detectable from RV variations. However, only on the super-luminous RGB ( 0.1 Hz, or 100 R/R⊙ 200), the RV amplitude of stellar oscillations becomes substantial enough to reach or exceed the RV threshold of 1.4 km/s (Katz et al., 2019) for faint stars of the Gaia RVS instrument ( 12 mag). As this regime is not probed by the range in APOKASC 3, possible false positive binary solutions produced by intrinsic stellar RV variations do not bias this comparison. We therefore assumed the completeness constant throughout the studied evolutionary states.
@@ 3.3 Evolution of the binary fraction for 2RC giants
The progenitors of secondary clump stars are early F- and A-type stars, which exhibit a different type of pulsations ( Dor and Scuti, Aerts et al., 2010, and references therein). Therefore, these are not represented in the MS sample of Mathur et al. (2022). We measure the binary fraction for the non-degenerate evolution starting from the low-luminosity RGB.
Table 5 and Fig. 2 show that such relatively massive stars represent only a small fraction of the full APOKASC 3 sample. Consequently, their population and binary counts suffer from small-sample statistics, and all differences found between the subsamples were deemed to be insignificant. The only marginally significant difference that we found from the numbers in Table 6 was that the binary fraction of the full sample of non-degenerate RGB stars (9.6%) is 17% richer in binaries than the full sample of degenerate RGB (8.2%). This increased richness aligns with the differences in the binary fraction of their MS progenitors, as observed in large binary statistics (Moe & Di Stefano, 2017; Offner et al., 2023).
4 Coevolution of the primary mass and the orbit
The key parameter determining the evolution of a star is its initial mass (Serenelli et al., 2021, and references therein). It defines the stellar structure and timescales of the evolution. In a binary system, the total mass of the components in solar units (), together with the semimajor axis, (in AU) or orbital period in years, defines the system’s configuration through the third law of Kepler (1619),
@@ 4 Coevolution of the primary mass and the orbit
Tidal interactions gain importance when the radius of one of the stellar components grows to a significant fraction of the semimajor axis. The redistribution of angular momentum and dissipation of kinetic and potential energy from tides into heat drives the changes in the orbits and the spins. This process leads the system toward a state of minimum energy, where the orbit is circularized, and the stellar spins become aligned and synchronized (Zahn, 1977; Ogilvie & Lin, 2007; Mathis, 2015). Tidal interaction occurs in waves as a dynamical tide or as a hydrostatic adjustment in the equilibrium tide, whereby the stellar structure, which depends primarily on the stellar mass, determines the dominant tidal mechanism. The various manifestations of tides are intensively studied through the distributions of orbital parameters in large samples from open clusters or from the Gaia mission, providing valuable constraints on tidal dissipation and orbital evolution (e.g., Mirouh et al., 2023; Dewberry & Wu, 2025).
Mass loss during stellar evolution, such as stellar winds, also affects orbital evolution by removing mass and angular momentum from the system (Miglio et al., 2012). Following Kepler’s third law (Eq. 1), this leads to an increase in the orbital period (Zahn & Bouchet, 1989; Soberman et al., 1997) and also to the spin-down of the mass-loosing star (e.g., Skumanich, 1972).
@@
If the radius of one of the binary components exceeds the orbit of the companion, the system will enter the short-lived, most likely destructive common-envelope (CE) phase, which would end the red-giant phase of the star abruptly with the CE ejection, exposing the “naked” He-core as a white dwarf (Ivanova et al., 2013). Due to the hydrodynamical complexity and their short timescales, this phase is difficult to model and poorly understood (Han et al., 2002). However, several parameters are identified to play a significant role in determining the fate of the binary system. The primary determining factor is the semimajor axis relative to the maximum stellar radius of the components, as this governs both tidal interaction strength and the Roche-lobe radius. Furthermore, the mass ratio between the two stars, which influences the difference in their evolutionary timescales (e.g., Miglio et al., 2014; Beck et al., 2018a; Grossmann et al., 2025)
@@
Stellar ages are typically inferred from model fitting, based on single-star evolutionary models assuming steady mass loss (e.g., Pinsonneault et al., 2025). Episodic or interaction-induced mass loss in binary systems will significantly over- or underestimate the ages of the post-interaction stellar products.
@@ 4.1 Mass distributions separated by evolutionary state
To isolate evolutionary effects in the binary sample, we constructed reference mass distributions for the entire sample, separated by evolutionary state. Since the mass distributions differ between the APOKASC 3 and APO-K2 catalogs, we analyze these two catalogs separately to avoid potential systematic biases in the samples skewing the result. The resulting KDE, shown in Fig. 4, shows that the mass distribution among the full samples is similar. Furthermore, Fig. 4 presents KDEs for the mass distributions for samples selected from various binary indicators. In the RGB phase, the mass distribution of all binary markers closely resembles that of the full sample, with only minor deviations.
@@
We suggest that the upward shift of the median mass in the binary sample is caused by the removal of lower-mass systems that arise from disruptive and merging events along the RGB, as shorter-period binaries undergo a CE phase, mostly leading to merger or envelope stripping of the RGB primary, before such stars can reach the tip of the giant branch and continue into the RC phase. This mechanism preferentially removes systems with and days and introduces a stronger mass dependence on the binary fraction than for the MS progenitors. Because the orbital period is the main parameter determining whether a system undergoes the CE phase, we expect a mass–period relation in the He-core-burning phase.
@@ 4.2 Mass dependency of the orbital parameters
For intermediate periods (8 Porb [d] 104), orbital eccentricities of MS binaries follow a Maxwellian “thermal” distribution (Zahn, 1989; Raghavan et al., 2010; Moe & Di Stefano, 2017), indicating that the probability density increases linearly with , i.e., systems with higher eccentricities are more common than low or circularized eccentricities which is still present in the eccentricities of MS stars in Fig. 5.
The combined large sample of binary stars with orbital solutions from the TBO with seismically inferred evolutionary states and masses from the APOKASC 3 and APO-K2 allows us to populate the eccentricity-period (e-P) plane (Fig. 5) and study the binary evolution as a function of stellar evolution (Fig. 6). We find a significant reduction in the ranges of orbital period and eccentricity from the MS into the He-core-burning phase. The short-periodic end of the period distribution shifts from a few days for MS stars to about 20 days for RGB. For stars in the He-core burning phase, the short periodic end is about 200 and 500 days for 2RC and RC stars. This picture is consistent with the findings of Dewberry & Wu (2025), who quantified comparable trends from a large-sample study of low-mass stellar binaries in the Gaia DR3 catalog.
@@ 4.3 Post-common envelope systems
We find numerous short periodic systems among the systems shown in the e-P plane in Fig. 5. Such systems are most likely artifacts of longer-periodic systems in the TBO. While red giants could exist in short-period systems, it is unlikely that such primaries would exhibit solar-like oscillations due to the effects of tidal interactions. We therefore excluded systems with Porb,TBO 10 days from the analysis but show them in Fig. 5.
@@
Only 7.5% of all giants with enhanced surface rotation are found to be in binary systems. Although this is slightly above the overall binary rate for RC stars, it is notable, as Patton et al. (2024) reported, based on all available binary indicators from Gaia DR3 and APOGEE RVs, a binary fraction of 732% for the entire sample of giants with enhanced surface rotation. This is significantly higher than the expected binary fraction of low-mass MS stars, indicating that these are the products of mergers and disruptive events.